OpenStudy (anonymous):

Express the complex number in trigonometric form. -2i

OpenStudy (anonymous):

you need two numbers \(r\) and \(\theta\) \[r=|2i|\] which should be pretty obvious what do you get ?

OpenStudy (anonymous):

2(cos(3Pi/2)+i sin(3Pi/2))------> expand and get the answer

OpenStudy (anonymous):

if it is not clear let me know but if you guess you will no doubt be right

OpenStudy (anonymous):

@satellite73 Do you mean i as in an imaginary number..? I have to use this for vectors and the demoivre theorem so wouldn't the "i" stand as an "i" in a vector?

OpenStudy (anonymous):

hold the phone lets go slow

OpenStudy (anonymous):

\(i\) in this case is the "imaginary unit" meaning \(i=\sqrt{-1}\) or \(i^2=-1\)

OpenStudy (anonymous):

the number \(-2i\) lives in the complex plane it is here 

OpenStudy (anonymous):

to write in trig form as \[-2i=r\left(\cos(\theta)+i\sin(\theta)\right)\] you need two numbers, \(r\) and \(\theta\)

OpenStudy (anonymous):

\(r\) is the absolute value the distance from zero what is that number?

OpenStudy (anonymous):

2? in this case?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

in general the absolute value of \(a+bi\) is \(\sqrt{a^2+b^2}\) but in this case it is completely obvious that is is \(2\)

OpenStudy (anonymous):

now you need \(\theta\)

OpenStudy (anonymous):



OpenStudy (anonymous):

that angle what is it?

OpenStudy (anonymous):

3pi/2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is also \(-\frac{\pi}{2}\) and an infinite number of other angles, it is not unique, but \(\frac{3\pi}{2}\) will work

OpenStudy (anonymous):

that means \[-2i=2\left (\cos(\frac{3\pi}{2})+i\sin(\frac{3\pi}{2})\right)\][

OpenStudy (anonymous):

ohhhh... now I get it :):) thank you so muuchh

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@satellite73 sorry but if it were to be 6-6i what would happen? since it's getting subtracted from 6?